Gold Sale. At 25% off, a jeweler is selling a gold chain for $465. What was the list price. I have been trying to resolve it and I am getting $581. When I try to do 25% of it I am not getting back the $465. Can you please help?
Thanks

Answer by askmemath(368)

At 25% off, the Selling price is only 75%
Which means that 75%of CostPRice= 465
What will 100% equal??
(100x465)/75 = $620
Check:
25% off means 25/100 off 0r 1/4th off
1/4th of 620 = 1/4 x 620 = 155
620-155 = 465 Yay!!!